OpenStudy (anonymous):

Explain how to find the zeros of the function f(x) = 2x3 – 9x + 3. Help, please:)

OpenStudy (anonymous):

f(x)=2x^3-9x+3

OpenStudy (dumbsearch2):

Let x = (a+b). Then x^3 = (a+b)^3 = a^3 + 3 a b (a + b) + b^3 = a^3 + 3 a b x+ b^3, or x^3 - 3 a b x - (a^3+b^3) = 0 Solve 3 a b = 9, and - (a^3+b^3) = 3 Then the answer is x = a + b x1= 3^(2/3)/(1/2 (-1 + i sqrt(11)))^(1/3) + (3/2 (-1 + i sqrt(11)))^(1/3) x2 = -(((3/2)^(2/3) (1 +i sqrt(3)))/(-1 +i sqrt(11)^(1/3)) - 1/2 (1 -i sqrt(3)) (3/2 (-1 +i sqrt(11)))^(1/3) x3 = -(((3/2)^(2/3) (1 -i sqrt(3)))/(-1 +i sqrt(11))^(1/3)) - 1/2 (1 +i sqrt(3)) (3/2 (-1 + i sqrt(11)))^(1/3)

OpenStudy (dumbsearch2):

Does that help, a little? :3

OpenStudy (anonymous):

It's a bit confusing but I'll find a way. Thanks you:)

OpenStudy (anonymous):

If anyone else has different methods of solving something like this please share:)

OpenStudy (mathteacher1729):

@jackoo two questions: 1. Are you sure you copied the problem correctly (I tried a few variations of it, none have "nice" roots. 2. What sort of answer are you looking for? Deriving the cubic equation? Numerical approximation? Graphing?

OpenStudy (dumbsearch2):

write it out in factored form, then mulitply out. say you want the zeros to be -2, 3 and 1/2 the write out (x+2)(x−3)(2x−1)

OpenStudy (dumbsearch2):

that obviously has zeros at -2, 3 and 1/2. multiply out to put it in standard form. i get 2x^3−3x^2−11x+6 but you can make up your own

OpenStudy (dumbsearch2):

just pick the zeros, write in factored form, and then multiply out. much easier than last problem, yes.

OpenStudy (anonymous):

Yes, I wrote it incorrectly but I corrected the equation. f(x)=2x^3-9x+3

OpenStudy (dumbsearch2):

1) write in factored form (don't show anyone) 2) multiply it out 3) factor again, you will get what you started with 4) now you know the zeros

OpenStudy (anonymous):

Thank you @dumbsearch2 and @mathteacher1729 do you have a different method?

OpenStudy (dumbsearch2):

(BTW if it helps, I would greatly appreciate if you could mark my answer as correct. ^_^)

OpenStudy (anonymous):

Will do:)

OpenStudy (dumbsearch2):

A lot of questions here have been asked previously.